OpenStudy (anonymous):

Write and then solve for y = f(x) the differential equation for the statement: "The rate of change of y with respect to x is inversely proportional to y^4."

OpenStudy (welshfella):

dy/dx = k / y^4

OpenStudy (anonymous):

Wouldn't it be x / y^4?

OpenStudy (anonymous):

no ,it is inversely proportional y^4 only.

OpenStudy (anonymous):

So how would I solve it?

OpenStudy (anonymous):

\[\sqrt[5]{5x^2/2+C}\]?

OpenStudy (anonymous):

But with a 'k' for 'x'?

OpenStudy (anonymous):

y^4dy=kdx integrate both sides \[\frac{ y^5 }{ 5 }=kx+c\]